# Thickness dependent rare earth segregation in magnetron deposited NdCo4.6 thin films studied by Xray reflectivity and Hard Xray photoemission

### v4

@@ Abstract
Magnetic anisotropy in disordered rare-earth-transition metals (RE-TM) compounds arises from RE atoms occupying asymmetric environments within the TM lattice. However, the underlying mechanism that promotes such environments remains not fully understood. In this study, we investigate amorphous NdCo4.6 thin films deposited by magnetron sputtering, where the magnetic anisotropy evolves with thickness from in-plane to out-of-plane orientation above 40 nm. X-ray reflectivity measurements revealed the progressive formation of an additional layer between the 3 nm Si capping layer and the NdCo film with increasing film thickness. To probe the composition and distribution of Co and Nd near the surface, Hard X-ray Photoemission Spectroscopy (HAXPES) was performed on films ranging from 5 nm to 65 nm in thickness using incident photon energies of 7, 19, and 13 keV. These correspond to inelastic electron mean free paths of 7.2-12.3 nm in cobalt. The cobalt atomic concentration, deduced from HAXPES at the Nd 3d and Co 2p excitations, was consistently below the nominal value and varied with both thickness and photon energy. This indicates segregation of RE atoms at the film surface, which becomes more pronounced with increasing thickness. Background analysis of the Co 2p and Nd 3d peaks supports this conclusion. The thickness of the Nd segregated layer is estimated in 2-3 nm. Nd segregation provides evidence for Nd diffusion toward the surface during thin film growth, what demonstrates that incorporating neodymium into the cobalt lattice incurs an energy cost which is associated with strain due to the volume mismatch between the two elements. Reducing this strain energy promotes anisotropic atomic environments for those Nd trapped into the cobalt lattice, thereby explaining the emergence of perpendicular anisotropy and its dependence on film thickness in these RE-TM compounds.
@@ I INTRODUCTION
Amorphous NdCo4.6 is a ferromagnetic intermetallic RE-TM compound with modest perpendicular magnetic anisotropy (PMA) at room temperature (RT), whose anisotropy energy quality factor is less than 1. This property makes it less anisotropy dominant and more flexible for implementation in various applications. They are produced at RT and can be stacked or combined with other materials giving rise to devices with interesting functionalities. Most of them are based on striped magnetic domains formed in thin films which have interesting properties such as a well-defined constant period, reconfigurable orientation, rotatable anisotropy [1], and topological textured features[2]. This compound is used, for instance, in bidimensional array composites with tunable exchange bias[3], in reconfigurable magnonic devices [4, 5] and in domain wall racetracks [6].
The PMA of RE-TM compounds is caused by the strong spin-orbit coupled 4f electron of the RE atoms. This orbital is nearly isolated from the valence atoms. Magnetic coupling between the spin of the TM and the 4f electrons occurs owing to the intratomic exchange with their valence 6s and 5d electrons which are spin polarized by the 3d TM electrons [7]. This coupling is weak, in the range of meV, and competes in energy with the crystal field at the RE[8, 7, 9] which fixes the orientation of its 4f angular moment that defines the anisotropy of the compound. The asymmetry of the RE atomic environment is, therefore, key for understanding the PMA properties of RE-TM alloys[10, 11, 12]. Structural asymmetry has been demonstrated in other RE-TM compounds such as TbFe by difraction[12], transmission electron microscopy[13] and extended Xray absorption fine spectroscopy (EXAFS)[14, 15].
Understanding how RE enters this type of asymmetric environment is an old problem owing to the complexity of the system[16, 17, 18]. However, the most accepted explanation for thin films deposited using magnetron sputtering methods, which have incident atom energies above those of thermal evaporation [19], is based on lowering the surface energy during thin film growth[20, 21]. Following this explanation, a majority of RE are in asymmetric environments because the corresponding adatom locations during growth are those that reduce the surface energy the most, creating a structural texture that gives rise to the resulting anisotropic environments. Thus, to control the PMA and the magnetic properties of the NdCo compounds it is neccesary to elucidate the parameters that condition the adatom diffusion during thin film growth.
@@ II EXPERIMENT
The samples were co-deposited using three confocal magnetron sputtering guns, Co, Nd and Si. The Co gun was normal to the substrate whereas the Nd and Si guns formed an angle of 30∘ on opposite sides to the normal of the sample. The Ar pressure during deposition was 3 mbar, with a base pressure of 1 mbar. The thin films were deposited on the native oxide of boron doped silicon substrates at a rate of 2 Åsec. The voltage bias and power applied to the magnetron guns during deposition were approximately 430 V and 0.10 kW for Co, and 330 V and 0.03 kW for Nd. All the films had a protective capping layer of 3 nm thick silicon. Deposition rates and concentration were calibrated using a quartz balanced and XRR.
The atomic concentration of the NdCo compound, NdCo4.6, was chosen because it gives the highest magnetic moment and PMA constant . Pure cobalt thin films of the same thickness were deposited in the same set of experiments to obtain a reference for cobalt magnetizations and layer interfaces. The magnetization and hysteresis loops (HL) of the samples were measured using a vibrating sample magnetometer (VSM) and an alternating gradient magnetometer (AGM). XRR measurements were performed in a two-axis difractometer (PANalytical X’pert) adapted to reflectometry measurements using a photon energy of 8.04 keV (Cu ).
@@ III.1 Magnetic Characterization
Figure 1 shows the in-plane field dependent magnetization (HLs) for pure cobalt and NdCo4.6 thin films of similar thickness obtained by magnetometry at RT. The shape of the HLs of cobalt was independent of the film thickness. This is shown in figure 1(a) where the HLs of the cobalt thin films with thicknesses ranging from 5 to 50 nm are superimposed. Their coercive field is 20 Oe, which is practically the same as that measured in the NdCo4.6 thinner films (fig. 1(b)). Owing to the normal incidence of Co atoms, they have negligible in-plane magnetic anisotropy. Their magnetic remanence was close to 60.
@@
The HLs measured in the NdCo4.6 samples with thicknesses below 40 nm are typical of soft magnets with in-plane anisotropy. Figure 1(b) shows them superimposed because they barely change with thickness. Their magnetic easy axis (EA) is perpendicular to the plane of incidence of the Nd and Si atoms, confirming the oblique incidence origin of their in-plane anisotropy[42, 43, 44]. Their EA HLs have almost 100 remanence, indicating that PMA energy is small compared with the in-plane anisotropy at these thicknesses. The hard axis (HA) has a saturation field, , of approximately 200 Oe at 30 nm, increasing to almost 400 Oe at lower thicknesses. The opening of the hysteresis loop in the HA might indicate a possible out of plane anisotropy component. As the thickness of the samples increased above 40 nm, the increased, the remanence decreased and the in-plane increased, as shown in figure 1(c). These are all signs of increasing magnetic anisotropy in the alloy. The HL of the 65 nm thick sample, in figure 1(d), shows a which is 10 times the measured in the samples with thickness lower than 40 nm. In addition, there is almost no difference between the HL measured along the in-plane EA and the HA: the in-plane anisotropy becomes overwhelmed by the out-of-plane anisotropy.
@@
The magnetization of both cobalt and NdCo4.6 thin films is approximately 18 lower at 5 nm, but reaches its saturation value for thicknesses above 20 nm, indicating cobalt interdiffusion at the silicon interfaces. XRR measurements shows interfaces widths of the order of 4 to 5 Å. The dominant in-plane magnetic anisotropy at low thickness, where the magnetization is the lowest, proofs that the PMA energy at low thickness is smaller than at higher thickness. The magnetization of the cobalt thicker films was the expected for an atomic magnetic moment of Co of 1.7 . The magnetization of the thickest NdCo films was approximately 800 emu/cm3. The expected magnetization of the NdCo4.6 compound, , can be calculated from the atomic concentration, the neodymium and cobalt atomic volumes determined from their related densities (8.9 gr/cm3 and 6.9 gr/cm3), and their magnetic moments and . For the atomic concentration of NdCo4.6, . The expected magnetic moment of the 4f orbital in Nd is related to the electronic configuration . However, this value is difficult to reach at RT because of the weak exchange coupling between Co and Nd atoms, disorder and possible presence of Nd unbonded to Co [24, 25]. In addition, cobalt reduces its magnetic moment not only by bonding to neodymium but because of the reduced cobalt bond length registered by EXAFS [15, 45], which contributes to the reduction of the atomic magnetic moment of Nd. This was observed in previous XMCD measurements[24] in samples with similar atomic concentrations. By taking into account all these factors, it is estimated that the magnetic moment of Co and Nd in these samples is roughly 1.35 and 1.1 , respectively. As it will be shown from the HAXPES data, these relatively low magnetic moments can hardly be attributed to oxygen contamination.
@@
The anisotropy energy of the analyzed samples can be estimated from . PMA energy is 0.84 erg/cm3 in the 45 nm film and 2.4 erg/cm3 in the 65 nm thick film. The in-plane anisotropy energy of the samples with 30 nm thickness or lower is approximately 0.3 erg/cm3. The apparent sharp transition from in-plane to out-of-plane anisotropy, which happens in a thickness difference of only 10 nm, is explained because the magnetic anisotropy is measured with the applied magnetic field in the plane of the samples. The transition occurs for two reasons: the gradual increase in PMA energy with thickness and the increase in thickness itself. At a critical thickness, both quantities become large enough to allow the formation of stripe domains. After this transition, figure 1 shows that the PMA energy increases gradually with thickness. The analysis of thicker samples shows that the saturation value for the PMA energy is reached around 65 nm.
@@ III.2 Xray Reflectivity
Figure 2(c) presents two reflectivity curves of a 95 nm thick NdCo4.6 film, where the low-frequency amplitude modulation associated with the top double layer is more pronounced than in the 65 nm thick layer shown in (b). The fitting parameters for these top layers are essentially identical. Both reflectivity curves were measured on the same sample, with the X-ray plane of incidence oriented either parallel ( sign) or perpendicular ( sign) to the plane of incidence of the Nd and Si magnetron guns.
The differences in the amplitude of the Kiessig oscillations between the two orientations arise from variations in the roughness at the substrate interface. This asymmetry is also observed in thinner films, though the effect is less significant. It originates directly from the oblique incidence of Nd atoms during thin-film growth, which establishes a preferential diffusion direction perpendicular to the Nd incidence plane. As a result, the deposited tend to expand on the surface in the direction perpendicular to the plane of incidence creating a corrugation in the roughness of the interfaces which is anisotropic.
@@ III.3.1 Valence band region
Figure 3 compares the region near the valence band (VB) of NdCo4.6 films at two different thickness, 5 nm and 10 nm, with the VB of a pure cobalt thin film obtained with 7 keV Xray photons. The spectra were normalized to the intensity of the Co 3p peak at 59.5 eV. The spectrum of the 5 nm thick alloy has a signal related to silicon and oxygen excitations in the region between 6 eV ang 16 eV, corresponding to the VB of SiO [49]. This film has a much larger contribution from the silicon substrate than the other analyzed films because its thickness was lower than the of the photoemitted electrons for the 7 keV incident photon energy. The VB spectrum of the 10 nm thick alloy is representative of the NdCo4.6 thicker films because it is no different from the obtained in the 30 and 65 nm thin films. Therefore, the differences between the shown VB spectra indicates the regions where Si and O excitations are expected. The difference in shape between the Nd 5p peaks in the spectra of the NdCo compound is likely caused by the O 2s peak located at a binding energy (BE) of approximately 25 eV, whose more pronounced intensity comes from the native SiO interface of the substrate. The intensity of this O 2s peak is almost undetectable in the VB spectrum of the pure cobalt film.

### v5

@@ Abstract
Magnetic anisotropy in disordered rare-earth-transition metals (RE-TM) compounds arises from RE atoms occupying asymmetric environments within the TM lattice. However, the underlying mechanism that promotes such environments remains not fully understood. In this study, we investigate amorphous NdCo4.6 thin films deposited by magnetron sputtering, where the magnetic anisotropy evolves with thickness from in-plane to out-of-plane orientation above 40 nm. X-ray reflectivity measurements revealed the progressive formation of an additional layer between the 3 nm Si capping layer and the NdCo film with increasing film thickness. To probe the composition and distribution of Co and Nd near the surface, Hard X-ray Photoemission Spectroscopy (HAXPES) was performed on films ranging from 5 nm to 65 nm in thickness using incident photon energies of 7, 10, and 13 keV. These correspond to inelastic electron mean free paths of 7.2-12.3 nm in cobalt. The cobalt atomic concentration, deduced from HAXPES at the Nd 3d and Co 2p excitations, was consistently below the nominal value and varied with both thickness and photon energy. This indicates segregation of RE atoms at the film surface, which becomes more pronounced with increasing thickness. Background analysis of the Co 2p and Nd 3d peaks supports this conclusion. The thickness of the Nd segregated layer is estimated in 2-3 nm. Nd segregation provides evidence for Nd diffusion toward the surface during thin film growth, what demonstrates that incorporating neodymium into the cobalt lattice incurs an energy cost which is associated with strain due to the volume mismatch between the two elements. Reducing this strain energy promotes anisotropic atomic environments for those Nd trapped into the cobalt lattice, thereby explaining the emergence of perpendicular anisotropy and its dependence on film thickness in these RE-TM compounds.
@@ I INTRODUCTION
Amorphous NdCo4.6 is a ferromagnetic intermetallic RE-TM compound with modest perpendicular magnetic anisotropy (PMA) at room temperature (RT), whose anisotropy energy quality factor is less than 1, being the PMA energy and the total magnetization of the alloy. This property makes it less anisotropy dominant and more flexible for implementation in various applications. They are produced at RT and can be stacked or combined with other materials giving rise to devices with interesting functionalities. Most of them are based on striped magnetic domains formed in thin films which have interesting properties such as a well-defined constant period, reconfigurable orientation, rotatable anisotropy [1], and topological textured features[2]. This compound is used, for instance, in bidimensional array composites with tunable exchange bias[3], in reconfigurable magnonic devices [4, 5] and in domain wall racetracks [6].
The magnetic moment of a RE element originates from its strongly spin-orbit-coupled 4f orbital, which is largely isolated from the valence electrons. Magnetic coupling between the TM spin and the 4f electrons of the RE is facilitated by intratomic exchange interactions involving the 6s and 5d valence electrons, which are themselves spin-polarized by the 3d electrons of the TM [7]. This coupling is relatively weak, on the order of meV, and competes with the crystal field at the RE site, which determines the orientation of the 4f angular momentum [8, 7, 9]. This orientation defines the anisotropy of the Nd magnetic moments at each of their locations in the compound. The PMA observed in RE-TM compounds emerges when there is a preference in the symmetry axis of the RE crystal fields for the out-of-plane alignment of its magnetic moment. Understanding the mechanism that originates such a preferential asymmetry in the atomic environment of the RE within the alloy is therefore key for controlling the PMA properties of RE-TM alloys [10, 11, 12]. Structural asymmetry has been confirmed in other amorphous RE-TM systems, such as amorphous TbFe, through techniques including diffraction [12], transmission electron microscopy [13], and extended X-ray absorption fine spectroscopy (EXAFS) [14, 15].
Understanding how RE enters this type of preferential asymmetric environment is an old problem owing to the complexity of the system[16, 17, 18]. However, the most accepted explanation for thin films deposited using magnetron sputtering methods, which have incident atom energies above those of thermal evaporation [19], is based on lowering the surface energy during thin film growth[20, 21]. Following this explanation, a majority of RE are in specific asymmetric environments because the corresponding adatom locations during growth are those that reduce the surface energy the most, creating a structural texture that gives rise to the resulting anisotropic environments. Thus, to control the PMA and the magnetic properties of the NdCo compounds it is neccesary to elucidate the parameters that condition the adatom diffusion during thin film growth.
@@ II EXPERIMENT
The samples were co-deposited using three confocal magnetron sputtering guns, Co, Nd and Si. The Co gun was normal to the substrate whereas the Nd and Si guns formed an angle of 30∘ on opposite sides to the normal of the sample. The Ar pressure during deposition was 3 mbar, with a base pressure of 1 mbar. The thin films were deposited on the native oxide of boron doped silicon substrates at a rate of 2 Åsec. The voltage bias and power applied to the magnetron guns during deposition were approximately 430 V and 0.10 kW for Co, and 330 V and 0.03 kW for Nd. All the films had a protective capping layer of 3 nm thick silicon. Deposition rates and concentration were calibrated using a quartz balance and XRR.
The atomic concentration of the NdCo compound, NdCo4.6, was chosen because it gives the highest magnetic moment and PMA energy . Pure cobalt thin films of the same thickness were deposited in the same set of experiments to obtain a reference for cobalt magnetizations and layer interfaces. The magnetization and hysteresis loops (HL) of the samples were measured using a vibrating sample magnetometer (VSM) and an alternating gradient magnetometer (AGM). XRR measurements were performed in a two-axis difractometer (PANalytical X’pert) adapted to reflectometry measurements using a photon energy of 8.04 keV (Cu ).
@@ III.1 Magnetic Characterization
Figure 1 shows the in-plane field dependent magnetization (HLs) for pure cobalt and NdCo4.6 thin films of similar thickness obtained by magnetometry at RT. The shape of the HLs of cobalt was independent of the film thickness. This is shown in figure 1(a) where the HLs of the cobalt thin films with thicknesses ranging from 5 to 50 nm are superimposed. Their coercive field is 20 Oe, which is practically the same as that measured in the NdCo4.6 thinner films (fig. 1(b)). Owing to the normal incidence of Co atoms during thin film deposition, they have negligible in-plane magnetic anisotropy. Their magnetic remanence was close to 60.
@@
The HLs measured in the NdCo4.6 samples with thicknesses below 40 nm are typical of soft magnets with in-plane anisotropy. Figure 1(b) shows them superimposed because they barely change with thickness. Their in-plane magnetic easy axis (EA) is perpendicular to the plane of incidence of the Nd and Si atoms, confirming the oblique incidence origin of their in-plane anisotropy[42, 43, 44]. Their EA HLs have almost 100 remanence, indicating that PMA energy is small compared with the in-plane anisotropy at these thicknesses. The hard axis (HA) has a saturation field, , of approximately 200 Oe at 30 nm, increasing to almost 400 Oe at lower thicknesses. The opening of the hysteresis loop in the HA might indicate a possible out of plane anisotropy component. As the thickness of the samples increased above 40 nm, the increased, the remanence decreased and the in-plane increased, as shown in figure 1(c). These are all signs of increasing magnetic anisotropy in the alloy. The HL of the 65 nm thick sample, in figure 1(d), shows a which is 10 times the measured in the samples with thickness lower than 40 nm. In addition, there is almost no difference between the HL measured along the in-plane EA and the HA: the in-plane anisotropy becomes overwhelmed by the out-of-plane anisotropy.
@@
The magnetization of both cobalt and NdCo4.6 thin films is approximately 18 lower at 5 nm, but reaches its saturation value for thicknesses above 20 nm, indicating cobalt interdiffusion at the silicon interfaces. XRR measurements shows interfaces widths of the order of 4 to 5 Å. The dominant in-plane magnetic anisotropy at low thickness, where the magnetization is the lowest, proofs that its PMA energy is smaller than at higher thickness. The magnetization of the cobalt thicker films was the expected for an atomic magnetic moment of Co of 1.7 . The magnetization of the thickest NdCo films was approximately 800 emu/cm3. The expected magnetization of the NdCo4.6 compound, , can be calculated from the atomic concentration, the neodymium and cobalt atomic volumes determined from their related densities (8.9 gr/cm3 and 6.9 gr/cm3), and their magnetic moments and . For the atomic concentration of NdCo4.6, . The expected magnetic moment of the 4f orbital in Nd is related to the electronic configuration . However, this value is difficult to reach at RT because of the weak exchange coupling between Co and Nd atoms, disorder and possible presence of Nd unbonded to Co [24, 25]. In addition, cobalt reduces its magnetic moment not only by bonding to neodymium but because of the reduced cobalt bond length registered by EXAFS [15, 45], which contributes to the reduction of the atomic magnetic moment of Nd. This was observed in previous XMCD measurements[24] in samples with similar atomic concentrations. By taking into account all these factors, it is estimated that the magnetic moment of Co and Nd in these samples is roughly 1.35 and 1.1 , respectively. As it will be shown from the HAXPES data, these relatively low magnetic moments can hardly be attributed to oxygen contamination.
@@
The anisotropy energy of the analyzed samples can be estimated from . PMA energy () is 0.84 erg/cm3 in the 45 nm film and 2.4 erg/cm3 in the 65 nm thick film. The in-plane anisotropy energy of the samples with 30 nm thickness or lower is approximately 0.3 erg/cm3. The apparent sharp transition from in-plane to out-of-plane anisotropy, which happens in a thickness difference of only 10 nm, might be explained because the competition between the shape anisotropy, which forces an in-plane magnetization, and the PMA. As the thickness increases, a gradual increase in PMA energy might occur. However, because of the shape anisotropy, the magnetic film needs to be arranged in a stripe magnetic domain configuration. This happens at a critical thickness, which is the size of the stripe domains. This size depends on and the magnetic exchange in the alloy[46]. After this transition, figure 1 shows that the PMA energy increases gradually with thickness. The analysis of thicker samples shows that the saturation value for the PMA energy is reached around 65 nm.
@@ III.2 Xray Reflectivity
Figure 2(c) presents two reflectivity curves of a 95 nm thick NdCo4.6 film, where the low-frequency amplitude modulation associated with the top double layer is more pronounced than in the 65 nm thick layer shown in (b). The fitting parameters for these top layers are essentially identical. Both reflectivity curves were measured on the same sample, with the X-ray wave vector oriented either parallel ( sign) or perpendicular ( sign) to the plane of incidence of the Nd and Si magnetron guns.
The differences in the amplitude of the Kiessig oscillations between the two orientations arise from variations in the roughness at the substrate interface. This asymmetry is also observed in thinner films, though the effect is less significant. It originates directly from the oblique incidence of Nd atoms during thin-film growth, producing a shadow effect which establishes a preferential diffusion direction for Co perpendicular to the Nd incidence plane. As a result, the Co deposited tend to expand on the surface in the direction perpendicular to the plane of incidence creating a corrugation in the roughness of the interfaces which is anisotropic.
@@ III.3.1 Valence band region
Figure 3 compares the region near the valence band (VB) of NdCo4.6 films at two different thickness, 5 nm and 10 nm, with the VB of a pure cobalt thin film obtained with 7 keV Xray photons. The spectra were normalized to the intensity of the Co 3p peak at 59.5 eV. The spectrum of the 5 nm thick alloy has a signal related to silicon and oxygen excitations in the region between 6 eV ang 16 eV, corresponding to the VB of SiO [50]. This film has a much larger contribution from the silicon substrate than the other analyzed films because its thickness was lower than the of the photoemitted electrons for the 7 keV incident photon energy. The VB spectrum of the 10 nm thick alloy is representative of the NdCo4.6 thicker films because it is no different from the obtained in the 30 and 65 nm thin films. Therefore, the differences between the shown VB spectra indicates the regions where Si and O excitations are expected. The difference in shape between the Nd 5p peaks in the spectra of the NdCo4.6 compound is likely caused by the O 2s peak located at a binding energy (BE) of approximately 25 eV, whose more pronounced intensity comes from the native SiO interface of the substrate. The intensity of this O 2s peak is almost undetectable in the VB spectrum of the pure cobalt film.